The following data show the probability distribution of the number of speakers sold daily in an appliance shop during the past several months, X={0, 1, 2, 3, 4, 5} P(X)={.07, .18, .27, .19, .08, .21}, what is the probability that at most 4 speakers will be sold?


1
Expert's answer
2022-03-08T12:52:07-0500

"P(X\le4)=1-P(X=5)=1-0.21=0.79."